OpenStudy (tad1):

Find an identity for cos5x in terms of cosx.

OpenStudy (amistre64):

cos(x+x+x+x+x) :)

OpenStudy (tad1):

I don't think that's what they had in mind. lol

OpenStudy (amistre64):

:) start with cos(x + 4x) and turn it into: cos(x)sin(4x) - cos(4x)sin(x) and work it all down to from there, if i see it right; but that is a pain

OpenStudy (tad1):

I've tried that and several other attempts. Perhaps its just because I hate arithmetic. the text has16cosx^5 - 20cosx^3 +5cosx

OpenStudy (amistre64):

cos(x)sin(4x) - cos(4x)sin(x) cos(x)[sin(x)sin(3x) + cos(x)cos(3x)] - [cos(x)cos(3x)-sin(x)sin(3x)]sin(x) cos(x) - sin(x) sin(x)sin(3x) + cos(x)cos(3x) cos(x)cos(3x)-sin(x)sin(3x) ........................................................... cos(x) sin(x) +cos(x) sin(x)sin(2x) + cos(x)cos(2x) cos(x)cos(2x) - sin(x)sin(2x) - sin(x) cos(x) -sin(x) cos(x)cos(2x) - sin(x)sin(2x) sin(x)sin(2x) + cos(x)cos(2x) ............................................................... yeah, its a pain :)

OpenStudy (anonymous):

Can you use the triple and double angle formulas to break it up? I started, but my brain is hurting. Split it into cos(3x + 2x) = cos(3x)cos(2x) - sin(3x)sin(2x). I worked the first term using cos(3x)=cos^3(x) - 3sin^2(x)cos(x) and cos(2x)=cos^2(x)-sin^2(x) and got 8cos^5(x)-7cos^3(x)+3cos(x). Haven't worked the second term though...but it looks promising, no?

OpenStudy (tad1):

It does. I haven't come up with coefficients that look that hopeful in many attempts

OpenStudy (anonymous):

\[Cos[5 x] = 5 Cos[x] - 20 Cos[x]^3 + 16 Cos[x]^5 \]

OpenStudy (anonymous):

Ok...my algebra is, I'm sure off, but for me it reduced to -8cos^5x - 12cos^3x +cosx which is not what your book says. But I think the idea is sound...but my algebra is wonky.

OpenStudy (anonymous):

Got it!

OpenStudy (anonymous):

whew!

OpenStudy (tad1):

beautiful. thanks to both of you.

OpenStudy (anonymous):

no problem. that was an outstanding refresher for me. it took putting down the pen and typing it out to keep my terms in place. :)

OpenStudy (anonymous):

A Mathematica solution with comments is attached.

OpenStudy (tad1):

could you try again. All I got was a black box off screen. Thanks.

OpenStudy (anonymous):

This site does't allow tiff files to be attached. I'll have to covert a tiff screen capture to jpg.

OpenStudy (tad1):

I'd appreciate it. I'll give you a medal in advance.

OpenStudy (anonymous):

Thanks for the medal. A jpg is attached. Sorry for the delay. Had to bring up a virtualized version of Ubuntu Linux and then use the Terminal command, Convert.

OpenStudy (anonymous):

There was an error in the last comment. Equate Cos[x]^5 to .... should have been typed, Equate Cos[5x] to ... Correct pdf version is attached.